OpenStudy (anonymous):

f(x)=2+3 cos(π/10)(x−3) Find f(2.6) Please help I keep getting an answer thats not one of my choices!!!

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

A. -.06842 B. 2.6 C. 4.9763 D. 5

OpenStudy (anonymous):

are you getting -6.999973133?

OpenStudy (anonymous):

no im getting 0.858732

OpenStudy (anonymous):

how are you getting that? i am just trying to find the mistake

OpenStudy (anonymous):

I'm putting it into my calculator but then i put it into wolfram alpha and i got the same thing

OpenStudy (anonymous):

how are you putting it into your calculator?

zepdrix (zepdrix):

You're probably messing up your order of operations somewhere. Or perhaps you're in degree mode by mistake. Trying entering it like this: \[2+3\times \cos\left(\left(\pi\div10)(2.6-3\right)\right)\]

OpenStudy (anonymous):

wow i got 4.999 so 5 thanks:)

zepdrix (zepdrix):

Hmm as I said before, you're in degree mode. Fix that.

zepdrix (zepdrix):

You want to be in Radians*